# Motion of spinning particles around black hole in a dark matter halo

### v1

@@ Abstract
The motion of a rapidly rotating object in curved spacetime is affected by the spin-curvature force, an effect captured in the motion of spinning test particles. Recently, Cardoso et al. [Phys. Rev. D 105, L061501 (2022)] found an exact solution describing a black hole immersed in a Hernquist distribution of dark matter. In this work, we investigate the motion of spinning particles around this black hole. We use the Mathison-Papapetrou-Dixon equation and the Tulczyjew spin-supplementary condition to calculate the effective potential, four-momentum, and four-velocity of the spinning particle. The equatorial motion of spinning test particles and the properties of the innermost stable circular orbit are further studied. We find that the existence of dark matter halos can significantly change the orbital eccentricity, energy, and the innermost stable circular orbit parameters of spinning test particles. Compared to the Schwarzschild black hole, dark matter halos bring the innermost stable circular orbit of a spinning test particle closer to the event horizon. These results could help us understanding the properties of black holes in dark matter halos.
@@ I Introduction
It is a very effective method to study the nature of a black hole by the motion of test particles near them. Changes in the background spacetime will result in changes in the motion state of the test particles. These changes will be reflected in the black hole shadow, accretion disk and so on. In previous studies, the effect of dark matter halos on the shadow of black holes has been studied Xavier:2023exm . However, in addition to photons, massive spinning particles or spinning stars also orbit black holes. In particular, the inspirals of stellar mass compact objects on the supermassive black hole will be detected by the space-based Observatory Laser Interferometer space antenna Babak:2017tow . This will allow us to accurately test the properties of supermassive objects at the center of galaxies and their surroundings. In the subleading corrections of dynamics of these so-called extreme-mass-ratio inspirals, the spin of smaller companion plays a crucial role Huerta:2011zi ; Piovano:2021iwv . The effect of spin of smaller companion can be described by the motion of a spinning test particle around a massive compact object Warburton:2017sxk ; Skoupy:2021asz ; Mathews:2021rod ; Piovano:2020zin ; Drummond:2023wqc . Thus, considering the dynamics of non-zero spinning particles around a black hole may provide insight into the motion properties of stars orbiting supermassive black holes. For a spinning test particle, its motion no longer follows a geodesic due to the spin curvature force Wald:1972sz ; Hanson:1974qy . In this case, based on the pole-dipole approximation of the spinning test particle, we can describe the motion of a rotating test particle moving against a curved background through the Mathison-Papapetrou-Dixon (MPD) equation Mathisson:1937ms ; Papapetrou:1951gr ; Corinaldesi:1951gr ; Tulczyjew:1959tr ; Dixon:1964gr . Previous literature have investigated the properties of spinning test particles in different black hole or other backgrounds Suzuki:1998vy ; Han:2008zzf ; Jefremov:2015gza ; Harms:2016ctx ; LukesGerakopoulos:2017cru ; Zhang:2017nhl ; Mukherjee:2018zug ; Zhang:2018omr ; Zhang:2019oet ; Antoniou:2019awm ; Nucamendi:2019qsn ; Zhang:2020qam ; Semerak:2015dza ; Mukherjee:2018dmm ; Toshmatov:2019bda ; Benavides-Gallego:2021lqn ; Yang:2022jno ; Zhang:2022qzw ; Chen:2024sbc ; Witzany:2023bmq ; Liu:2024lda . In this paper, we focus on the motion of spinning test particles in the black hole background with dark matter halo, and study the influence of dark matter halo on the orbital eccentricity and velocity of spinning test particles, especially the innermost stable circular orbit (ISCO) properties.
@@
The remainder of this paper is organized as follows. In Sec. II, we review the solution and the properties for a Schwarzschild black hole surrounded by a dark matter halo. Based on this solution, in Sec. III, we use the MPD equation and the Tulczyjew spin-supplementary condition to obtain the effective potential, four-momentum, and four-velocity of the spinning test particle. The equatorial motion of spinning test particles and the properties of the ISCO are further studied. Finally, some conclusions and discussions are provided in Sec. IV.
@@ III.1 MPD Equation and Equatorial Motion
where $S^{\mu\nu}$ is the spin tensor associated with the particle’s internal angular momentum, $P^{\mu}$, $u^{\mu}\equiv dx^{\mu}/d\tau$, and $\lambda$ are the four-momentum, the four-velocity along the trajectory, and the affine parameter, respectively. A simple count of the degrees of freedom in variables $S^{\mu\nu}$, $P^{\mu}$, and $u^{\mu}$ in the equations of motion (11) and (12) yields 13, while we have only 10 equations available. This insufficiency in solving for all variables arises from the fact that the center of mass of a spinning body in relativity is observer-dependent. Consequently, an extra “spin-supplementary condition” must be imposed to obtain a deterministic system. Various choices for this condition exist in the literature. In the present work, we employ the Tulczyjew spin-supplementary condition Tulczyjew:1959tr , expressed as
@@
[…]
@@
[…]
@@ III.2 Orbits
×$\times\left(\bar{e}-\frac{-B-\sqrt{B^{2}-4AC}}{2A}\right),$
@@
$\;\;r_{a}/M_{\text{BH}}\;\;$ $\;\;\;\;\;\;\;\;r_{p}/M_{\text{BH}}\;\;\;\;\;\;\;$ $\;\;\;\;\;\;\;\;\epsilon\;\;\;\;\;\;\;$ $\;\;\;\;\;\;\;\;N\;\;\;\;\;\;\;$
@@
Table 1: The apastron $r_{a}$, periastron $r_{p}$, eccentricity $\epsilon$, and number of rotations $N$ associate to the orbits in Fig. 2.
@@
At the end of this subsection, we focus on the circular orbits of spinning test particles. Since the four-velocity $u^{\mu}$ of a spin particle is not necessarily parallel to the four-momentum $P^{\mu}$, it is possible that the worldline of the spinning test particle is spacelike, which is obviously not physical. Therefore, we use superluminal constraint (39) to check whether the circular orbital motion of spinning test particles is superluminal. By numerical scanning, the orbital classification of black holes with different compactness in parameter space $(s-l)$ is shown in Fig. 4, where $l=j-s$ is the orbit angular momentum. Similarly, $\bar{l}=\frac{s}{mM_{\text{BH}}}$ represents the dimensionless orbit angular momentum. The gray region represents that the spin particle has a time-like stable circular orbit, the blue region represents that the particle has no stable circular orbit, and the yellow region represents that the spin particle has only a space-like stable circular orbit. It can be seen that, even in the case of extremely compactness dark matter hole with large mass $M$, the diagram of parameter space does not change much from the case of the Schwarzschild black hole. This is because although the dark matter halo can affect the value of the effective potential, it does not have much effect on the overall shape of the effective potential.
@@
Figure 4: Properties of the circular orbits for the spinning test particle in the $(s-l)$ parameter space. The range of $\bar{s}$ and $\bar{l}$ is $(-8,8)$. In the gray region, the test particle can have a stable timelike circular orbit. In the yellow region, the test particle has only the unphysical spacelike circular orbits. In the blue region, the particle does not have circular orbits.
III.3 ISCO
@@
Next, we use the above three conditions (45), (46), (47), and the superluminal constraint (39) to solve the ISCO of the spinning test particle. In Figs. 6,  8, and  7, we plot the radius $r_{\text{ISCO}}$, orbital angular momentum $\bar{l}_{\text{ISCO}}$, and energy $\bar{e}_{\text{ISCO}}$ with different parameters. From Fig. 6, we can see that the ISCO radius $r_{\text{ISCO}}$, orbital angular momentum $\bar{l}_{\text{ISCO}}$, and energy $\bar{e}_{\text{ISCO}}$ of spinning test particles all decrease with the spin $s$. When the spin $s$ is large, the trajectory of particle becomes spacelike, as shown in the yellow region in Fig. 6.
@@ IV Conclusions
In this paper, we investigated the equatorial plane motion of a spinning test particle around the black hole immersed in a dark matter halo. Because of the spin-curvature coupling, the equation of particle motion is described by the MPD equation rather than the geodesic equation. Considering the simple case of parallel or anti-parallel spin momentum $s$ and orbital angular momentum $j$, we studied how dark matter halos affect the motion of spinning test particles using the MPD equations. We found that the dark matter halo can make the ISCO of the spin particle smaller, that is, the inner edge of the accretion disk of the black hole with the dark matter halo closer to the event horizon. This is consistent with the conclusion of the ISCO of massless particles around black holes with dark matter halos Xavier:2023exm .
@@
We began by reviewing the solution for a spherically symmetric black hole surrounded by dark matter halos. Based on this solution, we use the MPD equation and Tulczyjew spin-supplementary condition to obtain the effective potential and four-velocity of the spinning test particle moving along the equatorial plane. By using the effective potential and four velocities, we obtain the orbital characteristics of the spinning test particles, which is shown in Fig. 2 and Tab. 1. The results showed that the apastron $r_{a}$ and eccentricity $\epsilon$ increase with $\mathcal{C}$ while the periastron $r_{p}$ and the number of turns $N$ decrease with $\mathcal{C}$. This suggests that dark matter halos do affect the orbits of spinning test particles. In addition, Fig. 4 shows a parametric diagram indicating whether a spinning particle has a circular orbit in the $s-l$ parameter space. We found that compared with the Schwarzschild black hole, the difference in the parameter graph of black holes with dark matter halos is not obvious. This is due to the small influence of dark matter halos on the overall shape of the effective potential. Finally, we studied the ISCO of the spinning test particles. From Fig. 6, we found that the spin $s$ of a particle significantly affects the radius $r_{\text{ISCO}}$, orbital angular momentum $l_{\text{ISCO}}$, and energy $e_{\text{ISCO}}$ of the ISCO. However, when the spin of the particle is large, the trajectories of the particle may become spacelike. By investigating the influence of the compactness $\mathcal{C}$ on the ISCO of the black hole, we found that dark matter halos make the radius $r_{\text{ISCO}}$ of a spinning test particle smaller, and the more compact the dark matter halo, the smaller the radius. And dark matter halos of the same compactness but different masses have slightly different results, which is shown in Figs. 7 and  8. These results showed that dark matter halos do have an effect on the motion of spinning test particles around the black hole. Although dark matter halos in the real world are less compact, increasingly accurate astronomical observations may detect these faint signs in the near future.

### v2

@@ Abstract
The motion of a rapidly rotating object in curved spacetime is affected by the spin-curvature force, an effect captured in the motion of spinning test particles. Recently, Cardoso et al. [Phys. Rev. D 105, L061501 (2022)] found an exact solution describing a black hole immersed in a Hernquist distribution of dark matter. In this work, we investigate the motion of spinning particles around this black hole. We use the Mathison-Papapetrou-Dixon equation and the Tulczyjew spin-supplementary condition to calculate the effective potential, four-momentum, and four-velocity of the spinning particle. The equatorial motion of spinning test particles and the properties of the marginally bound orbits, innermost stable circular orbits, and periodic orbits are further studied. We find that the existence of dark matter halos can significantly change the orbital eccentricity, energy, and the marginally bound orbits, innermost stable circular orbits, and periodic orbits parameters of spinning test particles. Compared to the Schwarzschild black hole, dark matter halos bring the marginally bound orbit and innermost stable circular orbit of a spinning test particle closer to the event horizon. These results could help us understand the properties of black holes in dark matter halos.
@@ I Introduction
It is a very effective method to study the nature of a black hole by the motion of test particles near them. Changes in the background spacetime will result in changes in the motion state of the test particles. These changes will be reflected in the black hole shadow, accretion disk and so on. In previous studies, the effect of dark matter halos on the shadow of black holes has been studied Xavier:2023exm . However, in addition to photons, massive spinning particles or spinning stars also orbit black holes. In particular, the inspirals of stellar mass compact objects on the supermassive black hole will be detected by the space-based Observatory Laser Interferometer space antenna Babak:2017tow . This will allow us to accurately test the properties of supermassive objects at the center of galaxies and their surroundings. In the subleading corrections of dynamics of these so-called extreme-mass-ratio inspirals, the spin of smaller companion plays a crucial role Huerta:2011zi ; Piovano:2021iwv . The effect of spin of smaller companion can be described by the motion of a spinning test particle around a massive compact object Warburton:2017sxk ; Skoupy:2021asz ; Mathews:2021rod ; Piovano:2020zin ; Drummond:2023wqc . Thus, considering the dynamics of non-zero spinning particles around a black hole may provide insight into the motion properties of stars orbiting supermassive black holes. For a spinning test particle, its motion no longer follows a geodesic due to the spin curvature force Wald:1972sz ; Hanson:1974qy . In this case, based on the pole-dipole approximation of the spinning test particle, we can describe the motion of a rotating test particle moving against a curved background through the Mathison-Papapetrou-Dixon (MPD) equation Mathisson:1937ms ; Papapetrou:1951gr ; Corinaldesi:1951gr ; Tulczyjew:1959tr ; Dixon:1964gr . Previous literature have investigated the properties of spinning test particles in different black hole or other backgrounds Suzuki:1998vy ; Han:2008zzf ; Jefremov:2015gza ; Harms:2016ctx ; LukesGerakopoulos:2017cru ; Zhang:2017nhl ; Mukherjee:2018zug ; Zhang:2018omr ; Zhang:2019oet ; Antoniou:2019awm ; Nucamendi:2019qsn ; Zhang:2020qam ; Semerak:2015dza ; Mukherjee:2018dmm ; Toshmatov:2019bda ; Benavides-Gallego:2021lqn ; Yang:2022jno ; Zhang:2022qzw ; Chen:2024sbc ; Witzany:2023bmq ; Liu:2024lda . In this paper, we focus on the motion of spinning test particles in the black hole background with dark matter halo, and study the influence of dark matter halo on the orbital eccentricity and velocity of spinning test particles, especially the marginally bound orbits (MBO), the innermost stable circular orbit (ISCO), and the periodic orbit properties.
@@
The remainder of this paper is organized as follows. In Sec. II, we review the solution and the properties for a Schwarzschild black hole surrounded by a dark matter halo. Based on this solution, in Sec. III, we use the MPD equation and the Tulczyjew spin-supplementary condition to obtain the effective potential, four-momentum, and four-velocity of the spinning test particle. The equatorial motion of spinning test particles and the properties of the MBO, ISCO, and periodic orbit are further studied. Finally, some conclusions and discussions are provided in Sec. IV.
@@ III.1 MPD Equation and Equatorial Motion
where $S^{\mu\nu}$ is the spin tensor associated with the particle’s internal angular momentum, $P^{\mu}$, $u^{\mu}\equiv dx^{\mu}/d\tau$, and $\lambda$ are the four-momentum, the four-velocity along the trajectory, and the affine parameter, respectively. A simple count of the degrees of freedom in variables $S^{\mu\nu}$, $P^{\mu}$, and $u^{\mu}$ in the equations of motion (11) and (12) yields 13, while we have only 10 equations available. This insufficiency in solving for all variables arises from the fact that the center of mass of a spinning body in relativity is observer-dependent. Consequently, an extra “spin-supplementary condition” must be imposed to obtain a deterministic system. Various choices for this condition exist in the literature Costa:2014nta . Due to the Tulczyjew spin-supplementary condition’s Tulczyjew:1959tr ability to solve for the ISCO over a larger range of particle spins Harms:2016ctx ; LukesGerakopoulos:2017cru and its better agreement with the results of Hamiltonian dynamics Harms:2016ctx , we adopt the Tulczyjew spin-supplementary condition in this paper. The expression for the Tulczyjew spin supplementary condition is Tulczyjew:1959tr
@@
$P^{t}u^{r}-P^{r}\!$ $\!=\!$ $\!-\frac{M_{\text{BH}}}{r\psi}\frac{DP_{\phi}}{D\lambda}+\frac{M_{\text{BH}}P_{\phi}}{r\psi}\left(\frac{1}{r}+\frac{\psi^{\prime}}{\psi}\right)u^{r},$
$P^{t}u^{\phi}-P^{\phi}\!$ $\!=\!$ $\!\frac{M_{\text{BH}}}{r\psi}\frac{DP_{r}}{D\lambda}-\frac{M_{\text{BH}}P_{r}}{r\psi}\left(\frac{1}{r}+\frac{\psi^{\prime}}{\psi}\right)u^{r}.$ (28)
@@
$\!\!\!\!\!\!\!P^{t}u^{r}\!-\!P^{r}\!$ $\!\!=\!\!$ $\!\frac{M_{\text{BH}}}{2r\psi}R_{\phi\nu\alpha\beta}u^{\nu}S^{\alpha\beta}\!+\!\frac{M_{\text{BH}}P_{\phi}}{r\psi}\left(\frac{1}{r}\!+\!\frac{\psi^{\prime}}{\psi}\right)\!u^{r},$
$\!\!\!\!\!\!\!P^{\phi}\!-\!P^{t}u^{\phi}\!$ $\!\!=\!\!$ $\!\frac{M_{\text{BH}}}{2r\psi}R_{r\nu\alpha\beta}u^{\nu}S^{\alpha\beta}\!+\!\frac{M_{\text{BH}}P_{r}}{r\psi}\left(\frac{1}{r}\!+\!\frac{\psi^{\prime}}{\psi}\right)\!u^{r}.$
@@ III.2 Effective potential
$\times\left(\bar{e}-\frac{-B-\sqrt{B^{2}-4AC}}{2A}\right),$
@@
$r_{a}/M_{\text{BH}}\;\;$ $r_{p}/M_{\text{BH}}\;\;\;\;\;\;\;$ $\epsilon\;\;\;\;\;\;\;$ N\;\;\;\;\;\;\;\;N\;\;\;\;\;\;\;
@@
Table 1: The apastron $r_{a}$, periastron $r_{p}$, eccentricity $\epsilon$, and number of rotations NN associate to the orbits in Fig. 2.
@@
At the end of this subsection, we focus on the circular orbits of spinning test particles. Since the four-velocity $u^{\mu}$ of a spin particle is not necessarily parallel to the four-momentum $P^{\mu}$, it is possible that the worldline of the spinning test particle is spacelike, which is obviously not physical. Therefore, we use superluminal constraint (39) to check whether the circular orbital motion of spinning test particles is superluminal. By numerical scanning, the orbital classification of black holes with different compactness in parameter space (sl)(s-l) is shown in Fig. 4, where l=jsl=j-s is the orbit angular momentum. Similarly, $\bar{l}=\frac{l}{mM_{\text{BH}}}$ represents the dimensionless orbit angular momentum. The gray region represents that the spin particle has a time-like stable circular orbit, the blue region represents that the particle has no stable circular orbit, and the yellow region represents that the spin particle has only a space-like stable circular orbit. It can be seen that, even in the case of extremely compactness dark matter hole with large mass $M$, the diagram of parameter space does not change much from the case of the Schwarzschild black hole. This is because although the dark matter halo can affect the value of the effective potential, it does not have much effect on the overall shape of the effective potential.
@@
Figure 4: Properties of the circular orbits for the spinning test particle in the (sl)(s-l) parameter space. The range of $\bar{s}$ and $\bar{l}$ is (8,8)(-8,8). In the gray region, the test particle can have a stable timelike circular orbit. In the yellow region, the test particle has only the unphysical spacelike circular orbits. In the blue region, the particle does not have circular orbits.
III.3 Marginally stable orbit
For a spinning particle, the MBO represents the smallest possible circular bound orbit with the smallest radius and an energy of $\bar{e}=1$. Given an effective potential $V_{\rm eff}$ controlling spinning particle motion, MBO satisfies the conditions Wei:2019zdf
$V_{\rm eff}=1$ (45)
$dV_{\rm eff}/dr=0.$ (46)
The radius and orbital angular momentum of MBO can be obtained by solving these equations. The radius $r_{\rm MBO}$ and angular momentum $l_{\rm MBO}$ are obviously dependent on the dark matter halo compactness $\mathcal{C}$ and the particle spin $s$. Figure 5 shows the numerical results of $r_{\rm MBO}$ and $l_{\rm MBO}$ as functions of the particle spin parameter $s$ and the dark matter mass $M$. We observe that both $r_{\rm MBO}$ and $l_{\rm MBO}$ decrease with $s$. When $\mathcal{C}$ increases, $r_{\rm MBO}$ decreases and $l_{\rm MBO}$ increases.
(a)  $M=0$
(b)  $M=200M_{\text{BH}},\mathcal{C}=1/5$
(c)  $\bar{s}=-0.5$
(d)  $\bar{s}=0.5$
Figure 5: Plots of the radius $r_{\text{MBO}}$ and orbital angular momentum $l_{\text{MBO}}$ of the spinning test particle for the different parameters $\mathcal{C}$ and $s$. The light yellow areas indicate that the particle orbits are spacelike.
III.4 ISCO
@@
Next, we use the above three conditions (47), (48), (49), and the superluminal constraint (39) to solve the ISCO of the spinning test particle. In Figs. 7,  9, and  8, we plot the radius $r_{\text{ISCO}}$, orbital angular momentum $\bar{l}_{\text{ISCO}}$, and energy $\bar{e}_{\text{ISCO}}$ with different parameters. From Fig. 7, we can see that the ISCO radius $r_{\text{ISCO}}$, orbital angular momentum $\bar{l}_{\text{ISCO}}$, and energy $\bar{e}_{\text{ISCO}}$ of spinning test particles all decrease with the spin $s$. When the spin $s$ is large, the trajectory of particle becomes spacelike, as shown in the yellow region in Fig. 7.
@@
III.5 periodic orbits
After investigating the ISCOs and MBOs, we now shift our focus to periodic orbits near a black hole surrounded by dark matter halos. These orbits are characterized by their repetitive trajectories, which arise when the frequency ratio between the radial ($\omega_{r}$) and azimuthal ($\omega_{\phi}$) oscillations is rational. Although generic orbits may exhibit irrational frequency ratios, they can always be approximated by nearby periodic orbits with arbitrary precision.
In the spherically symmetric geometry under consideration, the motion is governed by radial $r$-motion and angular $\phi$-motion. The apsidal angle, $\Delta\phi$, is calculated by integrating the change in azimuthal angle over one radial oscillation, from periapsis ($r_{1}$) to apoapsis ($r_{2}$) and back Levin:2008mq
$\Delta\phi=\oint d\phi$ $=$ $2\int_{\phi_{1}}^{\phi_{2}}d\phi=2\int_{r_{1}}^{r_{2}}\frac{\dot{\phi}}{\dot{r}}dr.$ (50)
The factor of 2 accounts for the symmetrical path followed by the particle. The apsidal angle is significantly influenced by the particle’s energy, angular momentum, and the spacetime structure around the black hole, as indicated by the presence of functions $a(r)$ and $b(r)$. Consequently, black holes with different parameter values will exhibit varying apsidal angles. For the bound orbital we’re studying, the angular momentum $l$ is only between $l_{\text{ISCO}}$ and $l_{\text{MBO}}$, which is
$l_{\text{ISCO}}<l<l_{\text{MBO}},$ (51)
Once the orbital angular momentum and the energy of the particle are given, we can find the periodic orbit of the spin particle by the frequency ratio $q$. The frequency ratio $q$ is defined as the relation between $\omega_{r}$ and $\omega_{\phi}$, expressed in terms of three integers: zoom $z$, whirl $w$, and vertex $v$ Levin:2008mq
$q=\frac{\omega_{\phi}}{\omega_{r}}-1=\frac{\Delta\phi}{2\pi}-1=w+\frac{v}{z}.$ (52)
For irrational values of $q$, the particle’s trajectory traces a precessing orbit, characterized by a precession angle $w=\Delta\Phi-2\pi$. Conversely, rational values of $q$ correspond to periodic orbits, where the particle returns to its initial position after a finite time. To illustrate the behavior of $q$ more directly, we plot the relation between $q$ and the energy $e$ of spinning test particle for different parameters in Fig. 10. It can be seen that, the rational number $q$ increases with the $\bar{e}$. In Fig. 10(a), we take the black hole as Schwarzschild case but take different values of spin angular momentum $s$, and it can be seen that $s$ has a significant effect on the distribution of qeq-e. In Fig. 10(b), we fixed $\bar{s}=0.5$ and explored the effect of different dark matter halo masses $M$ on $q$. The results show that the closer the system is to the Schwarzschild case, the greater the maximum energy corresponding to $q$.
Now we study the orbits of the spin particles in specific periodic orbits. Figure 11 shows periodic orbits using different combinations of integers $(z,w,v)$, with the value of $z$ determining the number of blades in the orbit’s shape. Larger $z$ values correspond to more complex trajectories with larger blade profiles. From Fig. 11(a), we can see that when the dark matter halo parameters are $M=10$ and $a_{0}=1000$, the periodic orbits of the spin particles almost coincide with those in the Schwarzschild black hole case. But we believe that as the compactness $\mathcal{C}$ of the dark matter halo increases, the difference becomes apparent. Moreover, even if the periodic orbits of the spin particles around the dark matter halo black hole are very similar to those of the Schwarzschild black hole, because of the long time scale of the orbit and the radiation of gravitational waves, this may cause the initial small difference to become apparent later. Therefore, there is a good chance that future highly sensitive gravitational wave detectors or telescopes will detect this difference, which will further validate or constrain black hole models with dark matter halos.
(a)  $M=0$
(b)  $\bar{s}=0.5,a_{0}=1000M_{\text{BH}}$
Figure 10: Plots of the relations of the rational number $q$ and the energy $\bar{e}$ for different parameters $s$ and $\mathcal{C}$. The orbit angular momentum $l$ set to $l=\frac{l_{\text{MBO}}+l_{\text{ISCO}}}{2}$.
(a)  $\bar{e}=0.959889$
(b)  $\bar{e}=0.962253$
(c)  $\bar{e}=0.962015$
(d)  $\bar{e}=0.962281$
(e)  $\bar{e}=0.962152$
(f)  $\bar{e}=0.962283$
(g)  $\bar{e}=0.954769$
(h)  $\bar{e}=0.962192$
Figure 11: Periodic orbits for different values of $(z,w,v)$ around a black hole with dark matter halo. The value of parameters are $M=10M_{BH}$, $a0=1000M_{BH}$, $\bar{s}=0.5$, and $\bar{l}=4$.
@@ IV Conclusions
In this paper, we investigated the equatorial plane motion of a spinning test particle around the black hole immersed in a dark matter halo. Because of the spin-curvature coupling, the equation of particle motion is described by the MPD equation rather than the geodesic equation. Considering the simple case of parallel or anti-parallel spin momentum $s$ and orbital angular momentum $j$, we studied how dark matter halos affect the motion of spinning test particles using the MPD equations. We mainly studied three kinds of orbits, namely MBO, ISCO, and periodic orbit. We found that the dark matter halo can make the MBO and ISCO of the spin particle smaller, that is, the inner edge of the accretion disk of the black hole with the dark matter halo closer to the event horizon. This is consistent with the conclusion of the ISCO of massless particles around black holes with dark matter halos Cardoso:2021wlq . Our research contributes to a deeper understanding of the complex gravitational interactions between black holes and dark matter halos, providing promising possibilities for future observations of the effects of dark matter in strong gravitational fields.
@@
We began by reviewing the solution for a spherically symmetric black hole surrounded by dark matter halos. Based on this solution, we use the MPD equation and Tulczyjew spin-supplementary condition to obtain the effective potential and four-velocity of the spinning test particle moving along the equatorial plane. By using the effective potential and four velocities, we obtain the orbital characteristics of the spinning test particles, which is shown in Fig. 2 and Tab. 1. The results showed that the apastron $r_{a}$ and eccentricity $\epsilon$ increase with $\mathcal{C}$ while the periastron $r_{p}$ and the number of turns $N$ decrease with $\mathcal{C}$. This suggests that dark matter halos do affect the orbits of spinning test particles. In addition, Fig. 4 shows a parametric diagram indicating whether a spinning particle has a circular orbit in the sls-l parameter space. We found that compared with the Schwarzschild black hole, the difference in the parameter graph of black holes with dark matter halos is not obvious. This is due to the small influence of dark matter halos on the overall shape of the effective potential. We studied the MBO, ISCO, and periodic orbit of the spinning test particles. From Fig. 5, we found that the spin $s$ of a particle and the mass of dark matter halos significantly affect the radius $r_{\text{MBO}}$, orbital angular momentum $l_{\text{MBO}}$ of the spinning particle. The same effect is observed for the ISCO. However, when the spin of the particle is large, the trajectories of the particle may become spacelike. By investigating the influence of the compactness $\mathcal{C}$ on the ISCO of the black hole, we found that dark matter halos make the $r_{\text{ISCO}}$ of a spinning test particle smaller, and the more compact the dark matter halo, the smaller the radius. And dark matter halos of the same compactness but different masses have slightly different results, which is shown in Figs. 8 and 9. Finally, the study of periodic orbits of spin particles shows that spin particles have periodic orbital motions just like non-spin particles. The presence of dark matter would change the orbital characteristics of periodic orbits. These results showed that dark matter halos do have an effect on the motion of spinning test particles around the black hole. Although dark matter halos in the real world are less compact, increasingly accurate astronomical observations may detect these faint signs in the near future.